Vegas Company has the following unit costs: Variable manufacturing overhead $ 25 Direct materials 20 Direct labor 19 Fixed manufacturing overhead 12 Variable marketing and administrative 7 Vegas produced and sold 10,000 units. If the product sells for $100, what is the operating profit under full absorption costing?
Answer #1

Sales revenue = 100*10,000 = 1,000,000

Cost of Goods Sold = (25+20+19+12)*10,000 = 760,000

Gross profit = 1,000,000 - 760,000 = 240,000

Operating profit = 240,000 - (7*10,000)

= 170,000